OpenStudy (anonymous):

What is the equation of a parabola with a vertex at (0, 0) and a focus at (0, 6).

OpenStudy (anonymous):

did you try to make a sketch?

OpenStudy (anonymous):

yes but I don't know how to make the equation

OpenStudy (anonymous):

is the parabola going to be vertical or horizontal?

OpenStudy (anonymous):

vertical?

OpenStudy (anonymous):

good. So which of the following equations is used? (x-h)^2 = 4p(y-k) or (y-k)^2 = 4p(x-h) ?

OpenStudy (anonymous):

I think its the first one but I'm sure

OpenStudy (anonymous):

yes it's the first one. If the x is being squared, the it's vertical, and if the y is being squared, it's horizontal. So we have decided that we are going to use (x-h)^2 = 4p(y-k). All we need to do now is to find the values of h,k, and p yes?

OpenStudy (anonymous):

okay. can you tell me how to find them?

OpenStudy (anonymous):

What is (h,k) called?